OpenStudy (anonymous):

\[\lim_{x \rightarrow 1}\frac{\int\limits_{1}^{x}\ln (3y-3y ^{2}+y ^{3})dy}{(1-x)^{3}}\]

OpenStudy (anonymous):

l'hopital for this one

OpenStudy (anonymous):

take derivative top and bottom get \[\lim_{x\rightarrow 1}\frac{x^3-3x^2+3x}{-3(1-x)^2}\]

OpenStudy (anonymous):

wow that was wrong!

OpenStudy (ash2326):

start by differentiating numerator and denominator \[ \lim_{x->1} \frac{\frac{d}{dx}\int_{1}^{x} \ln(3y-y^2+y^3) dy}{\frac{d}{dx} (1-x)^3}\] we get \[ \lim_{x->1} \frac{ \ln(3x-x^2+x^3)-0}{-3(1-x)^2}\]

OpenStudy (anonymous):

\[\lim_{x\rightarrow 1}\frac{\ln(x^3-3x^2+3x)}{-3(1-x)^2}\]

OpenStudy (anonymous):

i forgot the log, sorry

OpenStudy (anonymous):

Um the differentiation of the numerator is valid due to FTC 2 right?

OpenStudy (anonymous):

yup. derivative of integral is integrand

OpenStudy (ash2326):

now on differentiating again \[ \lim_{x->1} \frac{\frac{3x^2-6x+3}{x^3-3x^2+3x}}{-6(1-x)}\]

OpenStudy (anonymous):

gonna hae to do it again though

OpenStudy (anonymous):

ash has it

OpenStudy (anonymous):

@ash x\rightarrow 1 \[x\rightarrow 1\]

OpenStudy (anonymous):

Thanks sat :)

OpenStudy (anonymous):

sat: x \to 1 also work ;)

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

\[x\to\]

OpenStudy (anonymous):

YEsss :D

OpenStudy (anonymous):

jeez and all this time. pluse i spell "rightarrow" wrong half the time

OpenStudy (ash2326):

Let's differentiate it for the last time \[\lim_{x\rightarrow1} \frac{\frac{(6x-6)(x^3-3x^2+3x)-(3x^2-6x+3)(3x^2-6x+3)}{(x^3-3x^2+3x)^2}}{+6}\]

OpenStudy (anonymous):

\[\leftrightarrow\]

OpenStudy (anonymous):

haha I was there too :P

OpenStudy (anonymous):

i will post this in latexpractice

OpenStudy (anonymous):

Sat, have you seen this http://openstudy.com/users/foolformath#/updates/4f2d5be5e4b0571e9cba67c0

OpenStudy (anonymous):

wow!

OpenStudy (ash2326):

If we substitute x=1 , numerator is zero and denominator is finite so the answer is 0

OpenStudy (anonymous):

ok thank you :)

OpenStudy (ash2326):

welcome :)